Xorno Company is considering an investment of tk. 30,00,000 in new project. The new project is expected to last 20 years. It will have tk.2.5 lac salvage value at the end of its useful life. Company management’s required rate of return is 20%. The expected net income is tk. 50000. Net annual cash flow is tk. 65000 and at 13% interest rate, its present value is values tk. 2,10,74,400. Make the decision of capital budgeting in the following format:

Techniques

Calculation

Accept/reject

Reasons

Cash payback period

NPV

IRR

ARR


1
Expert's answer
2021-09-01T11:37:47-0400

Solution:

1.). Cash payback period = "\frac{Initial \;Investment}{Annual\; Cashflow}"

Initial investment = 3,000,000

Annual cashflow = 65,000

Cash payback period = "\frac{3,000,000}{65,000} =" 46 years and 2 months

Decision = Reject

This is because the project takes a longer period to recoup its initial investment which is very risky for the company.

 

2.). NPV = "CF\times [\frac{1 - (1+r^{n} )}{r} ] - Initial \; Investment"


NPV = "65,000\times [\frac{1 - (1+0.2^{-20} )}{0.2} ] + (2.5\times0.026) - 3,000,000"


NPV = 316,523 – 3,000,000 = (2,683,477)

NPV = (2,683,477)


Decision = Reject

This is because NPV is negative, which means that the project in which the company is investing does not provide a positive return and so the project should be rejected.

 


3.). IRR = Is defined as the discount rate at which the NPV of a project becomes zero.

Formula: 0 = NPV ="\sum \frac{CF}{(1+IRR^{t} )} - Initial \;Investment"

IRR using excel formula = -7%

Decision = Reject

This is because IRR is less than the minimum rate of return and it is also negative, hence the project will result in negative returns.

 

4.). ARR = "\frac{Average \;Annual\; Income}{(Initial \;investment - Salvage\; value)} \times 100"


ARR = "\frac{50,000}{(3,000,000 - 2.5)} \times 100 = 1.67\%"


Decision = Reject

The ARR is very low which indicates that the project will not produce positive returns.